# End-To-End Training and Testing Gamification Framework to Learn Human Highway Driving

### v1

@@ II Related works
The approach of deep reinforcement learning has also been tested for the development for self-driving vehicles. In this case the agent (the vehicles in our case) gathers the environmental information and tries to maximize its rewards based on a defined policy [11, 12, 13]. In case of self-driving cars, the agent could be rewarded for smooth driving and avoiding colliding with other vehicles and following the traffic rules [14, 15]. The concept of reverse imitation learning has also been studied to train vehicles to perform self driving. Here they have used a path planner algorithm to generate a set of paths and then use inverse reinforcement learning to always select the path that gives agents the highest reward [16, 17, 18, 19]. In this scenario also it requires a simulated environment and large amounts of data for training and testing.
@@
Recently studies in navigation [24, 25] have demonstrated the potential of end-to-end mapping image to control. As the network is required to clone the behavior of the training, this is termed as behavior cloning. Instead of using a complicated robotics approach where the vehicle should perform actions like localization, state estimation and navigation [26, 27, 28, 29, 30], the training of autonomous vehicle can be simplified by using the process of behavior cloning, where the neural network learns how to drive the vehicle by observing how a human or a trained agent drives the vehicle [31, 32]. Although this approach is not the best case for solving the problem of autonomy, this approach can be used in environments like highways and parking zones. This approach is similar to how a human learns how to drive rather than a robotics approach [33, 34, 35, 36, 37].
The approach of deep learning has also been used for segmented tasks such as obstacle detection as mentioned in [38, 39]. Here they combined the use of deep learning and geometric modeling to detect obstacles in the self driving car’s environment. The primary detection is done by the deep learning part but it takes help from the cues coming from the geometric model based on a stereo camera. There has also been several studies that show the current issues and challenges faced while using deep learning to implement scene detection and scene perception [40, 41, 42]. The common theme in most of the issues faced by deep learning is the lack readily available data or high amount of resources required to generate that amount of data.
The evaluation of deep learning techniques to perform object detection in different driving conditions has been studied [43, 44, 45, 46], here the use of transfer learning has been discussed to have an edge while training the neural networks to detect objects in different environments. Similar to the results in our approach transfer learning has produced a significant improvement while training and inference of the networks. The use of virtual data to train a deep learning network and inference it in the real world data has been tested [47, 48, 49, 50]. The use of video games to collect data and augment existing data-sets has been studied here [51, 52]. This helped them increase the size of the data, which is crucial for the training performance and also collect data that is hard to find in real life.

### v2

@@ II Related works
The approach of deep reinforcement learning has also been tested for the development for self-driving vehicles. In this case the agent (the vehicles in our case) gathers the environmental information and tries to maximize its rewards based on a defined policy [11, 12, 13, 14]. In case of self-driving cars, the agent could be rewarded for smooth driving and avoiding colliding with other vehicles and following the traffic rules [15, 16, 17]. The concept of reverse imitation learning has also been studied to train vehicles to perform self driving. Here they have used a path planner algorithm to generate a set of paths and then use inverse reinforcement learning to always select the path that gives agents the highest reward [18, 19, 20, 21, 22]. In this scenario also it requires a simulated environment and large amounts of data for training and testing.
@@
Recently studies in navigation [27, 28] have demonstrated the potential of end-to-end mapping image to control. As the network is required to clone the behavior of the training, this is termed as behavior cloning. Instead of using a complicated robotics approach where the vehicle should perform actions like localization, state estimation and navigation [29, 30, 31, 32, 33, 34], the training of autonomous vehicle can be simplified by using the process of behavior cloning, where the neural network learns how to drive the vehicle by observing how a human or a trained agent drives the vehicle [35, 36]. Although this approach is not the best case for solving the problem of autonomy, this approach can be used in environments like highways and parking zones. This approach is similar to how a human learns how to drive rather than a robotics approach [37, 38, 39, 40, 41].
The approach of deep learning has also been used for segmented tasks such as obstacle detection as mentioned in [42, 43]. Here they combined the use of deep learning and geometric modeling to detect obstacles in the self driving car’s environment. The primary detection is done by the deep learning part but it takes help from the cues coming from the geometric model based on a stereo camera. There has also been several studies that show the current issues and challenges faced while using deep learning to implement scene detection and scene perception [44, 45, 46, 47]. The common theme in most of the issues faced by deep learning is the lack readily available data or high amount of resources required to generate that amount of data.
The evaluation of deep learning techniques to perform object detection in different driving conditions has been studied [48, 49, 50], here the use of transfer learning has been discussed to have an edge while training the neural networks to detect objects in different environments. Similar to the results in our approach transfer learning has produced a significant improvement while training and inference of the networks. The use of virtual data to train a deep learning network and inference it in the real world data has been tested [51, 52, 53, 54]. The use of video games to collect data and augment existing data-sets has been studied here [55, 56]. This helped them increase the size of the data, which is crucial for the training performance and also collect data that is hard to find in real life.